# Good Intentions, Risky Inventions: A Method for Assessing the Risks and Benefits of AI in Mobile and Wearable Uses

### v2

@@ 7. Conclusion
1 Enhancing security through facial recognition using high-performance cameras in smartphones Yes 16, 5, 8 Article 6, Amendment 711 - Biometric identification and categorization of natural persons
2 Monitoring and analyzing crowd behavior using mobile cameras Yes 16, 11, 5, 8, 3 Article 6, Amendment 711 - Biometric identification and categorization of natural persons
3 Identifying individuals in criminal investigations using smartphones and wearable watches Yes 16, 10, 5, 8, 3, 11 Article 6, Amendment 711 - Biometric identification and categorization of natural persons
4 Monitoring children’s location for safety using wristbands Yes 3, 11, 16, 5, 8 Amendment 38 - AI Systems for perceiving or exploiting vulnerabilities of children and people
5 Monitoring elderly health and activity using wristbands Yes 3, 11 Amendment 38 - AI Systems for perceiving or exploiting vulnerabilities of children and people
@@ 8 Sharing location for coordination through smartwatches No 11, 5, 3 N/A
9 Tracking emotions using mobile app for improving emotional well-being Yes 5, 3 Article 6, Amendment 712 Biometric identification and categorization of natural persons
@@
12 Enhancing mental health treatment using smartwatches Yes 8, 3 Article 6 - Biometric identification and categorization of natural persons
@@ 19 Tracking expenses and monitoring personal finances using a mobile app No 12, 5, 8, 4, 1 N/A
23 Providing personalized learning recommendations using a tablet Yes 4, 8 Amendment 715, 716 - AI systems for assigning natural persons to educational and vocational training
@@
26 Monitoring physical movements for workplace safety using smartwatches Yes 3, 8, 5, 16, 11 Article 6 - AI systems as safety component and for monitoring and evaluating performance
27 Tracking work hours and tasks during remote work using a mobile app Yes 12, 5, 8, 4, 3 Article 6 - AI systems as safety component and for monitoring and evaluating performance
28 Tracking travel patterns for public transportation efficiency using smart cards Yes 17, 13, 12, 11, 5, 9, 8 AI Article 6 - AI systems as safety component
29 Monitoring health status and symptoms using headbands Yes 10, 9, 8, 3 AI Amendment 721 - evaluate the eligibility of natural persons for public assistance benefits and services
30 Tracking eligibility criteria and application status for social benefits using a mobile app Yes 10, 8, 5, 3, 1 AI Amendment 721 - evaluate the eligibility of natural persons for public assistance benefits and services
@@ 31 Tracking user preferences for personalized online shopping experience using a mobile app No 12, 8 N/A
32 Customizing music and video recommendations using earbuds Yes 12, 8 Amendment 740 - AI systems for recommending user-generated content available on the platform
33 Tailoring news and information feeds using a mobile app Yes 5, 4, 8, 12 Amendment 740 - AI systems for recommending user-generated content available on the platform
@@ 34 Tracking user interactions for enhancing user engagement using a mobile app No 8, 12 N/A
35 Facilitating social connections using a mobile app Yes 3, 8 Amendment 740 - AI systems for recommending user-generated content available on the platform
36 Monitoring harmful content using a mobile app Yes 10, 12, 5, 4, 3 Amendment 740 - AI systems for recommending user-generated content available on the platform
@@ 40 Tracking user motion for enhancing user experience using VR headsets No 5, 8, 3, 4 N/A
41 Monitoring audience reactions for facilitating interactive performances using mobile cameras Yes 5, 8, 4 Amendment 52, 191 - AI for inferring emotions, thoughts, states of mind individuals or groups
@@ 42 Analyzing user preferences for personalized content recommendations using a mobile app No 5, 8, 4 N/A
43 Preventing unauthorized access through biometric authentication using mobile scanners Yes 16, 5, 8 Article 6, Amendment 711 Biometric identification and categorization of natural persons
@@ 60 Improving research data management using a mobile app No 17, 9, 8, 4 N/A
64 Enhancing surveillance capabilities using mobile cameras * 16, 5, 8, 11 Article 5, 6 - Remote biometric identification systems in publicly accessible spaces used by law enforcement
65 Improving crime prediction using mobile cameras Yes 16, 5, 8 Article 6 - AI Systems for predicting criminal offence based on profiling of natural persons
@@
70 Enhancing voter participation using a mobile app Yes 16, 4 Amendment 739 - AI Systems for influencing the outcome of an election or referendum or the voting behaviour
71 Improving election transparency using a mobile app Yes 16, 10, 8, 9 Amendment 739 - AI Systems for influencing the outcome of an election or referendum or the voting behaviour
@@ 75 Improving content accessibility using mobile screen readers No 16, 8, 4, 9 N/A
76 Assisting visually impaired individuals using smart canes Yes 11, 16, 4, 10, 8, 3 Amendment 38 - AI Systems for perceiving or exploiting vulnerabilities of children and people
77 Supporting individuals with hearing impairment using earbuds Yes 11, 16, 4, 10, 8, 3 Amendment 38 - AI Systems for perceiving or exploiting vulnerabilities of children and people
@@ 80 Optimizing renewable energy production using heat sensors No 13, 12, 17, 11, 15, 9, 7 N/A
85 Facilitating remote court proceedings using a mobile app Yes 17, 10, 16, 11, 8, 9 Amendment 738 – AI systems for assisting in judicial authority in researching and interpreting facts
@@
87 Enhancing public participation in democratic processes using smart kiosks Yes 17, 16, 10, 11, 5, 8, 4, 9 Amendment 738 – AI systems for assisting in judicial authority in researching and interpreting facts
88 Improving public service delivery using a mobile app Yes 16, 11, 5, 8, 4, 3, 9 Amendment 67 - AI for applying for or receiving public assistance benefits and services
@@ 93 Improving international negotiation processes using a mobile app No 16, 17, 10, 4 N/A
99 Coordinating rescue operations using a mobile app Yes 17, 13, 8, 9, 3, 11, 16 Amendment 724 - AI systems for dispatching of emergency first response services.
100 Improving aid distribution efficiency using motion sensors Yes 12, 17, 16, 11, 1, 2, 3, 13, 8, 9 Amendment 721 - AI systems for evaluating the eligibility of natural persons for public assistance
101 Monitoring health status of aid recipients using wristbands Yes 17, 3, 9, 8, 11 Amendment 721 - AI systems for evaluating the eligibility of natural persons for public assistance
@@
131 Enhancing accessibility for disabled individuals using smart glasses Yes 11, 10, 8, 3 Article 6 - AI Systems for access to essential private services and public services and benefits

### v3

@@ 7. Conclusion
1 Enhancing security through facial recognition using high-performance cameras in smartphones Yes 16, 5, 8 Article 6, Amendment 711 - biometric identification/categorization of natural persons
2 Monitoring and analyzing crowd behavior using mobile cameras Yes 16, 11, 5, 8, 3 Article 6, Amendment 711 - biometric identification/categorization of natural persons
3 Identifying individuals in criminal investigations using smartphones and wearable watches Yes 16, 10, 5, 8, 3, 11 Article 6, Amendment 711 - biometric identification/categorization of natural persons
4 Monitoring children’s location for safety using wristbands Yes 3, 11, 16, 5, 8 Amendment 38 - perceive or exploit vulnerabilities of children and people
5 Monitoring elderly health and activity using wristbands Yes 3, 11 Amendment 38 - perceive or exploit vulnerabilities of children and people
@@ 8 Sharing location for coordination through smartwatches No 11, 5, 3 N/A
9 Tracking emotions using mobile app for improving emotional well-being Yes 5, 3 Article 6, Amendment 712 - biometric identification/categorization of natural persons
@@
12 Enhancing mental health treatment using smartwatches Yes 8, 3 Article 6 - biometric identification/categorization of natural persons
@@ 19 Tracking expenses and monitoring personal finances using a mobile app No 12, 5, 8, 4, 1 N/A
23 Providing personalized learning recommendations using a tablet Yes 4, 8 Amendment 715, 716 - assign natural persons to educational and vocational training
@@
26 Monitoring physical movements for workplace safety using smartwatches Yes 3, 8, 5, 16, 11 Article 6 - AI systems as safety component and for monitoring/evaluating performance
27 Tracking work hours and tasks during remote work using a mobile app Yes 12, 5, 8, 4, 3 Article 6 - AI systems as safety component and for monitoring/evaluating performance
28 Tracking travel patterns for public transportation efficiency using smart cards Yes 17, 13, 12, 11, 5, 9, 8 Article 6 - AI systems as safety component
29 Monitoring health status and symptoms using headbands Yes 10, 9, 8, 3 Amendment 721 - evaluate natural persons’ eligibility for public benefits and services
30 Tracking eligibility criteria and application status for social benefits using a mobile app Yes 10, 8, 5, 3, 1 Amendment 721 - evaluate natural persons’ eligibility for public benefits and services
@@ 31 Tracking user preferences for personalized online shopping experience using a mobile app No 12, 8 N/A
32 Customizing music and video recommendations using earbuds Yes 12, 8 Amendment 740 - recommend user-generated content available on the platform
33 Tailoring news and information feeds using a mobile app Yes 5, 4, 8, 12 Amendment 740 - recommend user-generated content available on the platform
@@ 34 Tracking user interactions for enhancing user engagement using a mobile app No 8, 12 N/A
35 Facilitating social connections using a mobile app Yes 3, 8 Amendment 740 - recommend user-generated content available on the platform
36 Monitoring harmful content using a mobile app Yes 10, 12, 5, 4, 3 Amendment 740 - recommend user-generated content available on the platform
@@ 40 Tracking user motion for enhancing user experience using VR headsets No 5, 8, 3, 4 N/A
41 Monitoring audience reactions for facilitating interactive performances using mobile cameras Yes 5, 8, 4 Amendment 52, 191 - infer emotions, thoughts, states of mind individuals or groups
@@ 42 Analyzing user preferences for personalized content recommendations using a mobile app No 5, 8, 4 N/A
43 Preventing unauthorized access through biometric authentication using mobile scanners Yes 16, 5, 8 Article 6, Amendment 711 - biometric identification/categorization of natural persons
@@ 60 Improving research data management using a mobile app No 17, 9, 8, 4 N/A
64 Enhancing surveillance capabilities using mobile cameras * 16, 5, 8, 11 Article 5, 6 - remote biometric identification in public spaces used by law enforcement
65 Improving crime prediction using mobile cameras Yes 16, 5, 8 Article 6 - predict criminal offence based on profiling of natural persons
@@
70 Enhancing voter participation using a mobile app Yes 16, 4 Amendment 739 - influence the outcome of election/referendum/voting behaviour
71 Improving election transparency using a mobile app Yes 16, 10, 8, 9 Amendment 739 - influence the outcome of election/referendum/voting behaviour
@@ 75 Improving content accessibility using mobile screen readers No 16, 8, 4, 9 N/A
76 Assisting visually impaired individuals using smart canes Yes 11, 16, 4, 10, 8, 3 Amendment 38 - perceive or exploit vulnerabilities of children and people
77 Supporting individuals with hearing impairment using earbuds Yes 11, 16, 4, 10, 8, 3 Amendment 38 - perceive or exploit vulnerabilities of children and people
@@ 80 Optimizing renewable energy production using heat sensors No 13, 12, 17, 11, 15, 9, 7 N/A
85 Facilitating remote court proceedings using a mobile app Yes 17, 10, 16, 11, 8, 9 Amendment 738 – assist in judicial authority in researching and interpreting facts
@@
87 Enhancing public participation in democratic processes using smart kiosks Yes 17, 16, 10, 11, 5, 8, 4, 9 Amendment 738 – assist in judicial authority in researching and interpreting facts
88 Improving public service delivery using a mobile app Yes 16, 11, 5, 8, 4, 3, 9 Amendment 67 - apply for or receive public assistance benefits and services
@@ 93 Improving international negotiation processes using a mobile app No 16, 17, 10, 4 N/A
99 Coordinating rescue operations using a mobile app Yes 17, 13, 8, 9, 3, 11, 16 Amendment 724 - AI systems for dispatching of emergency first response services
100 Improving aid distribution efficiency using motion sensors Yes 12, 17, 16, 11, 1, 2, 3, 13, 8, 9 Amendment 721 - evaluate natural persons’ eligibility for public benefits and services
101 Monitoring health status of aid recipients using wristbands Yes 17, 3, 9, 8, 11 Amendment 721 - evaluate natural persons’ eligibility for public benefits and services
@@
131 Enhancing accessibility for disabled individuals using smart glasses Yes 11, 10, 8, 3 Article 6 - access to essential private services and public services and benefits